### v1

# Let LLMs Speak Embedding Languages: Generative Text Embeddings via Iterative Contrastive Refinement

Abstract
Existing large language model (LLM)-based embeddings typically adopt an encoder-only paradigm, treating LLMs as static feature extractors and overlooking their core generative strengths. We introduce GIRCSE (Generative Iterative Refinement for Contrastive Sentence Embeddings), a novel framework that leverages autoregressive generation to iteratively refine semantic representations. By producing sequences of soft tokens optimized under contrastive objective, GIRCSE captures latent concepts and implicit semantics that encoder-only methods often miss. To guide this process, we propose an Iterative Contrastive Refinement (ICR) objective that encourages each refinement step to yield better representations. Extensive experiments show that GIRCSE outperforms strong LLM-based embedding baselines on the MTEB benchmark and instruction-following tasks. Moreover, GIRCSE exhibits an emergent test-time scaling property: generating more tokens at inference steadily improves embedding quality. Our results establish generative iterative refinement as a new paradigm for representation learning.
1 Introduction
Text embeddings are fundamental to a wide range of natural language processing (NLP) applications, including information retrieval, semantic search, clustering, and recommendation Karpukhin et al. (2020); Liu et al. (2024); Xiong et al. (2021). With the rise of large language models (LLMs), representation learning has advanced further: fine-tuning LLMs on large corpora now yields superior performance on several embedding benchmarks Tao et al. (2024).
However, current LLM-based embeddings typically adopt an encoder-style design: embeddings are extracted in a single forward pass with contrastive learning objectives, without leveraging the generative capacity of LLMs. This overlooks a key strength of pretrained LLMs—their ability to reason and iteratively refine through autoregressive generation Wei et al. (2022); Muennighoff et al. (2025). This raises a fundamental question: Can LLM-based embedding models also benefit from iterative generation? We hypothesize that generation enables iterative refinement of embeddings, allowing models to progressively consolidate semantics over multiple steps rather than encoding all semantics in a single pass.
Challenges. Designing effective generative embeddings presents several challenges. First, naive generation degrades embedding quality since pretrained LLMs are optimized for fluent text, not tokens aligned with semantic similarity (see Section 5.1). Second, unlike traditional language modeling, there is no clear generation target: it is unclear what content the model should generate to obtain universally useful embeddings. Third, existing embedding learning frameworks do not accommodate multi-step generative refinement. Therefore, it is necessary to develop new training paradigms that provide meaningful supervision for generative embeddings.
Figure 1: Top: Comparison between embedding LLMs that extract static representations and generative LLMs that can iteratively refine through reasoning. Bottom: Overview of GIRCSE. Our framework combines Soft Token Generation and Iterative Contrastive Refinement to enable end-to-end generative training.
Motivation. We argue that LLMs should learn to speak an embedding language: generating tokens not constrained by human readability but optimized for semantic representation. Crucially, these tokens should be discovered through end-to-end training jointly with contrastive objectives, enabling the model to generate semantically meaningful tokens for iterative embedding refinement.
Building on this motivation, we propose GIRCSEGenerative Iterative Refinement for Contrastive Sentence Embeddings—a novel framework that bridges this gap between generative LLM capabilities and embedding optimization. GIRCSE consists of two major innovations: (1) Soft Token Generation preserves differentiability for end-to-end contrastive training and captures richer semantics by retaining the diversity of the full probability distribution. (2) Iterative Contrastive Refinement (ICR) provides contrastive supervision at every generation step, forcing the early generated tokens to capture useful semantics while later tokens progressively refine representations. As illustrated in Fig. 1 and detailed in Section 5.2, this paradigm enables GIRCSE to generate instruction-aware refinement tokens (e.g.,”frustrated” and ”struggle”) that capture the implicit emotion beyond the surface text. In summary, we make the following contributions:
  • Novel embedding framework. We propose GIRCSE, a novel end-to-end framework that integrates autoregressive generation with contrastive learning. Unlike prior methods, GIRCSE generates soft refinement tokens without explicit targets, progressively distilling semantics into high-quality embeddings.
  • Superior performance. We compare GIRCSE with 18 state-of-the-art embedding models. By generating only up to 5–20 additional tokens, GIRCSE ranks within top 5–6 on MTEB and top 2–3 on instruction following, leading to the best overall ranking across benchmarks. Meanwhile, GIRCSE consistently shows stable improvements over reproduced fair baselines on different backbone and varying data scales.
  • Test-time scaling ability for text embedding. We demonstrate that GIRCSE exhibits consistent embedding quality improvements with increased refinement steps at inference time, representing a novel scaling paradigm for embedding models analogous to test-time compute scaling in reasoning LLMs.
2 Related Work
Early Embedding Models. Text embedding methods have evolved from traditional word-level representations to sophisticated neural approaches. Early methods like Word2Vec Mikolov et al. (2013) and GloVe Pennington et al. (2014) captured basic semantic relationships but lacked contextual understanding. The introduction of transformer-based models marked a significant advancement, with BERT-based approaches like Sentence-BERT Reimers & Gurevych (2019) and SimCSE Gao et al. (2021) establishing contrastive learning as the dominant paradigm for learning sentence representations.
LLM-based Embedding models. Recent works have successfully adapted LLMs for representation learning through various architectural and training modifications. E5-mistral Wang et al. (2024) is one of the early works that demonstrated that fine-tuning LLM could significantly outperform early stage methods. Recognizing that the unidirectional attention mechanism in LLMs may limit text embedding quality, LLM2Vec BehnamGhader et al. (2024) introduces a bidirectional attention mechanism combined with average pooling to enhance embedding quality. NV-Embed Lee et al. (2024) further improves the pooling method by incorporating an additional Latent Attention Layer and implements a two-stage training strategy to address the challenge of false negatives in non-retrieval tasks. BGE-en-icl Li et al. (2025) suggests that retaining the original framework of LLMs and leveraging in-context learning is the optimal approach for generating text embeddings.
Towards Generative Text Embedding. A smaller line of research has explored generative approaches for text embeddings. For example, Inbedder Peng et al. (2024) combines instruction finetuning with token generation, achieving strong performance on instruction-following tasks but showing limited generalization to broader tasks (see Table 2). As summarized in Table 1, most existing approaches differ only in pooling strategies or auxiliary training techniques, while generative embeddings remain largely underexplored.
Table 1: Comparison of LLM-based embedding models. ”Bidir.” indicates bidirectional attention, ”GP” means generated tokens pooling, ”TTS” refers to test-time scaling capability.
Method Attention Pooling LoRA Generation TTS Gen. Token Training Obj.
E5-mistral Causal EOS N/A sup. CL
SFR-Embedding Causal EOS N/A sup. CL
BGE-en-ICL Causal EOS N/A ICL & sup. CL
GRITLM Bidir. Avg. N/A NTP & sup. CL
LLM2Vec Bidir. Avg. N/A MLM & sup. CL
GTE-Qwen2 Bidir. Avg. N/A unsup. CL & sup. CL
NV-Embed-v1 Bidir. LAT N/A Two-stage sup. CL
Inbedder Causal GP Hard Token Instruction Tuning
GIRCSE Causal GP Soft Token sup. CL & ICR
3 GIRCSE: From Discriminative to Generative Embedding
We now detail our proposed generative embedding framework. Section 3.1 first establishes our core autoregressive embedding generation process, introducing the fundamental concepts and notation. Section 3.2 then details the soft token generation mechanism that enables differentiable optimization within this framework. Finally, Section 3.3 presents our iterative contrastive refinement objective, guiding the model towards progressively higher-quality representations.
3.1 Generative Embedding Framework
We consider a pretrained language model with parameters $\psi=\{\mathbf{E},\theta,\phi\}$, where $\mathbf{E}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the token embedding matrix, $\theta$ denotes the parameters of the Transformer decoder, and $\phi$ corresponds to the parameters of the LM head for next token generation. Here $d$ is the embedding dimension and $|\mathcal{V}|$ is the vocabulary size. Given an input sequence of $N$ tokens $\mathbf{T}=\{t_{1},t_{2},\ldots,t_{N}\}$ from vocabulary $\mathcal{V}$, we first obtain token embeddings $\mathbf{X}$ as:
$\mathbf{X}=(\mathbf{E}[t_{1}],\mathbf{E}[t_{2}],\ldots,\mathbf{E}[t_{N}])\in\mathbb{R}^{N\times d}.$ (1)
Next, our goal is to autoregressively generate a sequence of $K$ auxiliary soft tokens $\mathbf{S}=(\mathbf{s}_{1},\mathbf{s}_{2},\ldots,\mathbf{s}_{K})\in\mathbb{R}^{K\times|\mathcal{V}|}$ that iteratively refines the representation space. Each soft token $\mathbf{s}_{k}$ is generated autoregressively conditioned on the input sequence and previously generated tokens:
$p_{\psi}(\mathbf{S}\mid\mathbf{T})=\prod_{k=1}^{K}p_{\psi}(\mathbf{s}_{k}\mid\mathbf{T},\mathbf{S}_{<k}),$ (2)
where $p_{\psi}$ represents the generative distribution and $\mathbf{S}_{<k}=(\mathbf{s}_{1},\ldots,\mathbf{s}_{k-1})$ are the previously generated soft tokens. The soft tokens are then mapped into embedding space [^1] We defer the detailed soft token generation mechanism to Section 3.2, while here we focus on the overall framework, producing $\mathbf{D}=(\mathbf{d}_{1},\ldots,\mathbf{d}_{k})\in\mathbb{R}^{K\times d}$, and is concatenated with input embeddings $\mathbf{X}$ to further feed into the Transformer encoder $f_{\theta}$:
$\mathbf{H}=f_{\theta}([\mathbf{X};\mathbf{D}])=(\mathbf{h}_{1}^{(L)},\mathbf{h}_{2}^{(L)},\ldots,\mathbf{h}_{N+K}^{(L)})\in\mathbb{R}^{(N+K)\times d},$ (3)
where $\mathbf{h}_{i}^{(L)}$ denotes the hidden state of the $i$-th token at the final (i.e., $L$-th) layer. We then extract the representations corresponding to the generated soft tokens, and aggregate them into a single representation $\mathbf{z}$ via a pooling operation:
$\mathbf{z}=\mathcal{P}(\mathbf{G})=\frac{1}{K}\sum_{i=1}^{K}\mathbf{g}_{i},\;\;\mathbf{G}=(\mathbf{g}_{1},\ldots,\mathbf{g}_{K})=(\mathbf{h}_{N+1}^{(L)},\ldots,\mathbf{h}_{N+K}^{(L)})\in\mathbb{R}^{K\times d},$ (4)
where $\mathcal{P}$ denotes a general pooling function, with mean pooling as our default choice.
[…]
3.2 Soft Token Generation
A critical challenge in our autoregressive framework is to maintain the differentiability throughout the generation process during training. Traditional discrete token sampling would break gradient flow, preventing end-to-end optimization. We address this through a novel soft token generation mechanism that preserves continuous optimization while capturing rich semantic information.
At each generation step $k\in\{1,\ldots,K\}$, the generative distribution $p_{\psi}$ is instantiated via the LM head $\phi$. Let $\mathbf{h}^{\prime}_{k-1}=\mathbf{h}^{(L)}_{N+k-1}$ denote the last layer hidden representation produced by the encoder given the input sequence and the previously generated soft tokens up to step k1k-1, the LM head then produces a soft token $\mathbf{s}_{k}\in\mathbb{R}^{|\mathcal{V}|}$ as a probability distribution over the vocabulary:
$\mathbf{s}_{k}=\text{softmax}(\mathbf{W}_{\phi}\mathbf{h}^{\prime}_{k-1}+\mathbf{b}_{\phi}),$ (5)
where $\mathbf{W_{\phi}}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is the LM head weight matrix and $\mathbf{b}_{\phi}$ is the bias term. Given the soft token $\mathbf{s}_{k}$, its embedding $\mathbf{d}_{k}\in\mathbb{R}^{d}$ is obtained by computing a convex combination of all token embeddings according to their predicted probabilities:
$\mathbf{d}_{k}=\sum_{i=1}^{|\mathcal{V}|}s_{k,i}\mathbf{e}_{i},$ (6)
where $s_{k,i}$ is the $i$-th component of $\mathbf{s}_{k}$ and $\mathbf{e}_{i}$ is the $i$-th row of the embedding matrix $\mathbf{E}$. This soft token generation approach offers two advantages: (1) Differentiability: The weighted combination preserves gradients throughout the generation process, enabling end-to-end training with contrastive objectives. (2) Semantic Richness: Rather than collapse the next-token distribution into a single token, soft tokens capture the semantic diversity of the full probability distribution.
3.3 Guiding Generative Embedding with Iterative Contrastive Refinement
To guide the generative embedding process towards high-quality representations, we introduce an iterative contrastive refinement (ICR) objective that encourages each generation step to yield increasingly refined representations. ICR combines (1) Stepwise Contrastive Loss, which supervises each generation step with contrastive loss, and (2) Iterative Refinement Regularization, which enforces progressive embedding quality improvement for each step.
Stepwise Contrastive Loss. In autoregressive soft token generation, supervising only the final embedding (i.e., $K$-th generation step) might collapse intermediate steps into trivial or noisy representations. We instead apply contrastive supervision at every generation step. Concretely, for step $k$, we pool the first $k$ generated tokens to form an intermediate embedding $\mathbf{z}_{k}=\mathcal{P}(\mathbf{G}_{1:k})$ following Eq. 4. Given a query–document pair $(q,d^{+})$, we compute the contrastive loss for all generation steps as:
$\mathcal{L}_{\text{contrast}}=\sum_{k=1}^{K}\mathcal{L}_{k},\quad\mathcal{L}_{k}=-\log\frac{\exp\!\big(\sigma(\mathbf{z}_{k}^{q},\mathbf{z}_{k}^{d^{+}})/\tau\big)}{\sum_{d\in\mathcal{B}}\exp\!\big(\sigma(\mathbf{z}_{k}^{q},\mathbf{z}_{k}^{d})/\tau\big)},$ (7)
where $\mathcal{B}$ denotes the document set (both positive and negative documents), $\sigma$ is the cosine similarity function, and $\tau$ is the temperature hyperparameter. This stepwise supervision ensures all intermediate representations align with the contrastive objective, preventing early steps from drifting and providing richer supervision.
Iterative Refinement Regularization. We empirically observe that simply increasing the number of generation steps does not guarantee improved embedding quality, as LLMs often produce highly similar tokens which leads to redundant information in the multi-step process. To address this, we introduce a regularization term that encourages monotonic improvement across generation steps:
$\mathcal{L}_{\text{reg}}=\frac{1}{K-1}\sum_{k=1}^{K-1}\max\!\big(\log\mathcal{L}_{k+1}-\log\mathcal{L}_{k},\,0\big).$ (8)
This regularization term penalizes cases where later generation steps fail to outperform earlier ones. Finally, the overall fine-tuning objective for generative embeddings combines the two terms: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{contrast}}+\lambda\mathcal{L}_{\text{reg}}$, where $\lambda$ is a hyperparameter that balances contrastive alignment and refinement regularization.
4 Experiment
4.1 Experiment Setup
Backbone LLM. Following prior works Wang et al. (2024); Muennighoff et al. (2024), we adopt Mistral-7B Jiang et al. (2023) as the primary backbone and further validate it on Qwen2.5-7B Yang et al. (2024).
[…]
Evaluation. We evaluate on MTEB (English, v2) Enevoldsen et al. (2025), covering 41 datasets across 7 task types, reporting official leaderboard scores when available. Following InBedder Peng et al. (2024), we also evaluate on IntentEmotion and NYTClustering to test the instruction-following ability of embedding models. For more extensive comparison, we have also evaluated on TREC datasets used in FollowIR Weller et al. (2025) and BEIR Thakur et al. (2021), detailed results can be found in Appendix H.
Comparison Methods. We compare GIRCSE against four categories of text embedding models. (1) Non-LLM methods including encoder-based models such as E5-Large Wang et al. (2022), GTE-Large Li et al. (2023b), BGE-Large Li et al. (2023a), and UAE-Large Li & Li (2024). (2) LLM-based methods are instruction-tuned LLM embeddings, including LLM2Vec BehnamGhader et al. (2024), GritLM Muennighoff et al. (2024), E5-Mistral Wang et al. (2024), NV-Embed-v1 Lee et al. (2024), SFR-Embedding-2 Meng* et al. (2024), and gte-Qwen2 Li et al. (2023b). (3) Generative embeddings cover (i) two-stage approaches that expand text with an auxiliary LLM before re-encoding (see Appendix G for detail) and (ii) the end-to-end generative model Inbedder Peng et al. (2024). (4) Fair Baselines are included by re-implementing two paradigms on the same training data for fair comparison: (i) Causal-EOS (causal attention + EOS pooling) and (ii) Bidirectional-Avg (bidirectional attention + average pooling), equivalent to E5-Mistral and GritLM respectively but trained with less data.
4.2 Main Results
Table 2 reports the performance comparison across MTEB tasks and instruction-following benchmarks. We highlight the following observations:
Trade-off between generic tasks and instruction following. State-of-the-art non-generative embedding models achieve strong results on generic MTEB tasks but lag behind on instruction-following benchmarks. For example, gte-QWEN2 performs competitively on MTEB (rank 1) but drops notably on instruction-following tasks (rank 18). Similarly, E5-Mistral ranks 4 on MTEB but falls to 10 on instruction following. In contrast, generative embedding approaches such as Inbedder reverse this trend, achieving top instruction-following performance (rank 1), since it is explicitly trained for this setting, but performing poorly on MTEB (rank 20). A comparable trade-off is also observed in two-stage generative variants of non-generative models. For instance, E5-Mistral (w/ gen) improves on instruction following (rank 10 $\rightarrow$ 5) but degrades on MTEB (rank 4 $\rightarrow$ 12). Similar patterns are also observed for E5-Large (w/ gen) and GritLM (w/ gen).
GIRCSE overcomes trade-off and strikes a balanced performance. Unlike prior methods, GIRCSE delivers consistently strong results across both task categories. It not only outperforms fair baselines and competitive embedding models (e.g., GritLM, LLM2Vec), but also avoids the severe trade-offs observed in existing approaches. Specifically, GIRCSE ranks within the top 5–6 on MTEB and top 2–3 on instruction-following tasks, leading to the best overall rankings of 3.5 and 4.5 across benchmarks. Remarkably, while prior SOTA methods rely on multi-million–scale training datasets, GIRCSE achieves comparable or better performance with only 0.2M training examples. These results highlight GIRCSE as an efficient embedding model that achieves both strong general-purpose performance and robust instruction-following ability.
Table 2: Performance on MTEB and instruction-following tasks. Results obtained from the official MTEB leaderboard. Causal-EOS: causal attention with EOS pooling; Bidirectional-Avg: bidirectional attention with average pooling. Highlighted rows are our reproductions, trained on a smaller dataset (0.2M) for fair comparison. Bold = better than fair baselines with same backbone; * = statistically significant ($p<0.05$). For detailed performance of each MTEB dataset, please refer to Appendix I.
MTEB (English, v2) Instruct Following
Overall
Rank
Task Size Vol. Backbone Retr. Rerank. Clust. PairClass. Class. STS Summ. Avg. (Rank) IntEmo NYT Avg. (Rank)
# of datasets $\rightarrow$ - - - 10 2 8 3 8 9 1 41 1 1 2
Non-LLM Methods
E5-Large 0.3B 1B BERT 49.31 45.72 45.23 86.06 76.44 80.67 32.34 62.79 (18) 48.63 50.96 49.80 (14) 16.0
GTE-Large 0.3B 2B BERT 53.29 47.84 48.20 85.08 75.47 83.27 32.90 64.77 (14) 52.62 17.52 35.07 (18) 16.0
BGE-Large 0.3B 200M BERT 55.44 48.26 48.01 87.13 78.34 82.79 33.13 65.89 (13) 51.66 61.38 56.52 (8) 10.5
UAE-Large 0.3B 1M RoBERTa 55.91 48.35 47.86 87.25 79.08 84.37 30.13 66.40 (9) 50.49 60.54 55.52 (11) 10.0
LLM-based: Causal-EOS
E5-Mistral 7B 1.8M Mistral 57.62 49.78 51.44 88.42 79.85 84.32 36.57 67.97 (4) 48.84 65.06 56.95 (10) 7.0
SFR-Embedding-2 7B 1.7M Mistral 53.75 48.99 59.39 88.09 90.54 80.86 35.54 69.82 (2) 50.49 60.54 55.52 (11) 6.5
gte-Qwen2 7B 800M QWEN2 58.09 50.47 58.97 85.90 88.52 82.69 35.74 70.72 (1) 52.62 17.52 35.07 (18) 9.5
Fair Baseline 7B 0.2M Mistral 55.24 49.21 54.28 85.65 84.36 73.98 36.31 66.32 (10) 35.33 58.76 47.05 (15) 12.5
Fair Baseline 7B 0.2M QWEN2 51.10 47.49 55.26 84.46 80.10 74.71 33.21 64.18 (17) 66.14 14.71 40.42 (17) 17.0
LLM-based: Bidirectional-Avg
LLM2Vec 7B 1.5M Mistral 51.27 47.74 44.10 87.99 79.74 83.70 31.05 64.57 (15) 51.66 61.38 56.52 (8) 11.5
GritLM 7B 2M Mistral 54.95 49.59 50.82 87.29 81.25 83.03 35.65 67.07 (7) 39.30 79.25 59.28 (6) 6.5
NV-Embed-v1 7B 1.1M Mistral 60.13 49.16 49.50 87.05 84.11 82.20 31.40 68.32 (3) 52.61 60.62 56.62 (7) 5.0
Fair Baseline 7B 0.2M Mistral 55.41 48.74 54.57 86.34 84.94 75.87 36.09 66.96 (8) 21.45 66.42 43.94 (16) 12.0
Fair Baseline 7B 0.2M QWEN2 52.99 47.11 54.75 83.31 82.66 72.81 35.30 64.97 (16) 43.26 65.21 54.24 (13) 14.5
LLM-based: Two-Stage Generative Embedding
E5-Large (w/ gen.) 0.3B 1B BERT 45.06 43.87 45.37 81.02 72.70 77.35 31.59 59.85 (19) 51.34 51.67 51.51 (12) 15.5
E5-Mistral (w/ gen.) 7B 1.8M Mistral 57.20 49.18 53.02 84.26 75.97 79.52 31.73 65.92 (12) 58.64 60.89 59.77 (5) 8.5
GritLM (w/ gen.) 7B 2M Mistral 56.48 49.45 52.03 83.36 77.77 79.66 32.82 65.90 (11) 51.16 70.50 60.83 (4) 7.5
LLM-based: End2End Generative Embedding
Inbedder 7B 0.2M LLaMA2 12.50 39.21 51.24 61.17 72.41 74.41 17.24 50.32 (20) 89.68 64.65 77.17 (1) 10.5
GIRCSE 7B 0.2M Mistral 57.10 48.88 56.26 86.18 85.33 76.37 33.56 67.83* (5) 52.19 73.75 62.97 (2) 3.5
GIRCSE 7B 0.2M QWEN2 55.16 49.28 56.62 85.17 86.69 76.30 35.42 67.67* (6) 64.92 60.04 62.48 (3) 4.5
4.3 Ablation Study
To better understand the contributions of different components in GIRCSE, Table 3 represents an ablation study on generative embedding, stepwise loss (SL), and iterative refinement (IR). Starting from the variant without generation (i.e., the Causal-EOS baseline), we observe a substantial drop in performance across both MTEB and instruction-following tasks. Incorporating generative embedding alone yields consistent improvements across nearly all tasks. Adding SL provides further gains, particularly for classification and summarization, while the combination of SL and IR achieves the strongest overall performance. Overall, these results validate the effectiveness of our design in GIRCSE.
Table 3: Ablation study of GIRCSE with generative embedding (Gen.), stepwise loss (SL), and iterative refinement (IR). The variant without generation corresponds to the Causal-EOS baseline. Results are reported using the Mistral-7B backbone trained on 50K samples.
MTEB (English, v2) Instruct Following
Gen. SL IR Retr. Rerank. Clust. PairCls. Class. STS Summ. Avg. IntEmo NYT Avg.
50.55 48.97 49.91 85.27 80.36 75.76 34.02 63.84 33.11 58.76 47.05
53.17 48.32 52.74 84.75 78.34 78.70 33.86 65.21 48.00 64.93 56.47
54.97 48.86 52.07 85.04 78.63 78.87 35.28 65.69 53.88 66.37 60.13
55.53 48.26 53.71 84.87 79.53 78.93 34.19 66.27 62.70 73.75 62.97
5 Analysis on Generated Tokens
While Section 4.3 highlights the importance of generative embedding, it remains unclear how the generation process itself translates to the improved performance. To address this, we present a thorough analysis of the generation process. In Section 5.1, we first discuss how the embedding quality changes by varying the number of generated tokens at inference. Next, in Section 5.2, we conduct a qualitative analysis to understand what tokens are generated and how they evolve under different instructions. This analysis clarifies how iterative generation improves performance and what semantic signals are encoded in the embedding space.
Figure 2: Effect of generation length at inference. Top: GIRCSE consistently improves with longer generations (10–20 tokens) despite been trained on only 5 tokens. Bottom: Baseline models show degraded or fluctuated performance across generation lengths. Gray area indicates configurations beyond training length.
5.1 Effect of Generation Length at Inference
We first examine how performance varies with the number of generated tokens $K$ at inference. We evaluate $K\in\{1,3,5,10,15,20\}$ and compare against the non-generative baseline Causal-EOS. Results (Fig. 2) are reported on three LLM backbones and three representative MTEB tasks, with relative improvements measured against $K=1$ for GIRCSE and against the no-generation baseline for Causal-EOS. For a more comprehensive analysis, we further evaluate GIRCSE trained with two additional backbones: QWEN2.5-0.5B Yang et al. (2024) and LLaMA3.2-3B Dubey et al. (2024). We have the following two key findings:
(1) GIRCSE exhibits test-time scaling for embeddings. Increasing $K$ consistently improves performance across diverse tasks (e.g., STS, clustering, classification) and across model sizes. In contrast, the non-generative method (Causal-EOS) does not benefit from additional generation and often degrades in performance. This suggests that GIRCSE successfully learns an iterative refinement mechanism that converts additional inference computation into stronger semantic representations—analogous to test-time compute scaling in reasoning LLMs Muennighoff et al. (2025), but novel in the context of embedding models.
(2) ICR enables GIRCSE to generalize beyond training configurations. Although GIRCSE is trained with $K=5$, its performance improves monotonically within the training regime ($K=1,3,5$) and continues to improve even beyond it ($K=10,15,20$). This extrapolation capability suggests that our ICR training objective enables the learned refinement process to generalize beyond the training configuration, allowing GIRCSE to continue improving with additional inference steps. Overall, GIRCSE establishes test-time scaling as a new paradigm for embedding models, enabling controllable and training-free performance gains through adjustable generation length.
5.2 Qualitative Analysis on Generated Tokens
Having shown that generating more tokens improves performance, we next ask: what do these tokens capture? We analyze generations for the sentence “Why is it so hard to track down this card?” under two prompts: representing intention and emotion. At each generation step $k$, we collect the top-30 candidates from the soft token distribution $\mathbf{s}_{k}$, aggregate across steps, and report most frequent tokens in Table 4, alongside results from GIRCSE before contrastive fine-tuning. Before fine-tuning, GIRCSE (before FT) often yields generic or semantically weak tokens. After fine-tuning, we observe progressive semantic refinement that aligns with the results in Section 5.1. At early steps (1–5), GIRCSE generates core content words (e.g., why, hard, card). While at later steps, the outputs diverge by different instructions: intention produces tokens such as seek, elusive, inquiry, and emotion yields tokens like frustrating, struggle. This suggests multi-step generation acts as a semantic chain of thought, iteratively steering representations toward nuanced, instruction-aligned regions of the embedding space.
Table 4: Qualitative analysis of generated tokens. Gray indicates generic/stopword-like tokens. Yellow marks core input-related tokens shared across instructions. Instruction-specific expansions are shown in Green (intention) and Red (emotion).
Input Sentence “Why is it so hard to track down this card?”
Instruction “Represent the intention of this text.” “Represent the emotion of this text.”
GIRCSE (Before FT) this, so, do, how, i, is, it, the, we, what why, how, what, this, it, is, you, can
GIRCSE (Step 1–5) why, is, it, hard, track/tracking, card, this why, is, it, hard, track,tracking, card, this, so, difficult
GIRCSE (Step 6–10) [prev.] + seek, elusive, so [prev.] + frustrating, tough, persistent, struggle, challenging
GIRCSE (Step 11–20) [prev.] + question, inquiry [prev.] + perseverance, stuck, complicated
6 Discussion of Robustness and Learning Efficiency
To assess the robustness and learning efficiency of our method, we conduct comprehensive experiments across varying data scales and backbone architectures. Specifically, we train different models with {50K, 100K, 200K} training samples using three widely adopted open-source LLMs as base models: Qwen-0.5B, Llama-3B, and Mistral-7B. We compare GIRCSE against two fair baselines: Causal-EOS and Bidirectional-Avg. Fig. 3 shows that our method consistently outperforms both baselines across all data scales and model sizes. In particular, when trained with only 50K samples, our method improves over Causal-EOS by +5.7 points on Qwen-0.5B (61.2% vs. 55.5%) and by +2.8 points on Llama-3B (65.5% vs. 62.7%). Even with stronger backbones such as Mistral-7B, our approach still yields gains of +2.4 points (66.2% vs. 63.8%). The performance gap becomes more pronounced when the training data is limited. These findings indicate that our approach not only achieves superior performance across different scales of model size but also learns more effectively under limited training data.
Figure 3: Comparison of average MTEB scores (%) between GIRCSE and two fair baselines across three backbone LLMs and varying training sample sizes. GIRCSE consistently delivers superior performance, especially under limited-data settings.
7 Conclusion
We presented GIRCSE, a generative embedding framework that leverages autoregressive refinement to move beyond single-pass LLM encoders. By generating soft refinement tokens and training with iterative contrastive refinement, GIRCSE enables embeddings to progressively distill semantics rather than compressing them in one step. Experiments show that GIRCSE achieves state-of-the-art or competitive performance across benchmarks while introducing a novel scaling property: embedding quality improves with additional refinement steps at test time. These results highlight autoregressive generation as a powerful mechanism for embedding optimization and open new directions for scalable, semantically rich representations.
Appendix A Use-of-LLMs
In this work, we utilized large language models (LLMs) as part of the core research methodology. Specifically, we fine-tuned existing open-source LLMs (e.g., LLaMA-3 and Mistral) to develop embedding models. These pre-trained models served as the foundation for our experiments, and our main contributions build upon their architectures and representations. Additionally, an LLM-based assistant (OpenAI GPT-5) was used for minor writing support, including grammar checking and improving manuscript readability. All decisions regarding research design, fine-tuning strategies, experimental setup, and final interpretations were made solely by the authors.
Appendix B Reproducibility for Embedding Models
To facilitate reproducibility of our experiments, we provide links to all open-sourced embedding models used in this paper in Table 5. These links allow researchers to directly access the exact model checkpoints we relied on.
Table 5: List of models with links for reproducibility.
Appendix C Hyperparameter Settings
For both GIRCSE and the baseline models we re-implement for comparison, we use nearly identical fine-tuning hyperparameters across different model sizes (QWEN-0.5B, Llama-3B, and Mistral-7B) and training data scales (50k, 100k, and 200k examples).
We adopt Low-Rank Adaptation (LoRA) Hu et al. (2022) for efficient fine-tuning, setting the rank to 64 and the scaling factor $\alpha$ to 32. The default learning rate is $1\text{e-}5$ with a warmup ratio of 0.1. The only exception is Llama-3B, for which we use a learning rate of $1\text{e-}4$ to address convergence issues. For other hyperparameters, we set the temperature of the contrastive loss (Eq. 7) to 0.02 across all models, and the weighting coefficient $\lambda$ in GIRCSE for balancing contrastive alignment and refinement regularization to 1. Due to limited computational resources, we train with a batch size of 2 and accumulate gradients over 8 steps, resulting in an effective batch size of 16. All models are fine-tuned for a single epoch.
For GIRCSE, we set the number of generated tokens $K$ to 5 during training to avoid the high computational cost of multiple autoregressive forward passes. During inference, we increase $K$ to 20 to enable longer generations and improved embedding refinement, while mitigating the computational overhead using KV-cache techniques.
Appendix D Pseudo Code for GIRCSE
Algorithm 1 provides a pseudo-code summary of GIRCSE. We include this block to clarify the core computation steps and facilitate reproducibility. The algorithm specifies how input tokens are embedded, how soft embeddings are autoregressively generated across multiple steps, and how the final representation is obtained through pooling. While implementation details may vary, this summary highlights the essential components needed to reproduce our method.
Input: Input tokens $(t_{1},\ldots,t_{N})$, generation steps $K$, embedding matrix $\mathbf{E}$, LLM decoder $f_{\theta}$, LM head parameters $W_{\phi}$ and $b_{\phi}$, pooling function $\mathcal{P}$
Output: Final embedding $\mathbf{z}\in\mathbb{R}^{d}$
1ex
$\mathbf{X}\leftarrow(\mathbf{E}[t_{1}],\ldots,\mathbf{E}[t_{N}])$
// Embed the input tokens
$\mathbf{D}\leftarrow[\,]$
// Initialize the list of generated embeddings
for $k=1$ to $K$ do
 $\mathbf{H}\leftarrow f_{\theta}([\mathbf{X};\mathbf{D}])$
 // Forward with input and generated tokens
 $\mathbf{h}^{\prime}_{k-1}\leftarrow\mathbf{H}[N+k-1]$
 // Last hidden state for next-token prediction
 $\mathbf{s}_{k}\leftarrow\mathrm{softmax}(\mathbf{W}_{\phi}\mathbf{h}^{\prime}_{k-1}+\mathbf{b}_{\phi})$
 // Compute soft token distribution
 $\mathbf{d}_{k}\leftarrow\sum_{i=1}^{|\mathcal{V}|}s_{k,i}\mathbf{e}_{i}$
 // Compute soft embedding
 $\mathbf{D}\leftarrow\mathbf{D}\|\mathbf{d}_{k}$
 // Append soft embedding for next step
$\mathbf{H}\leftarrow f_{\theta}([\mathbf{X};\mathbf{D}])$
$\mathbf{G}\leftarrow(\mathbf{H}[N+1],\ldots,\mathbf{H}[N+K])$
// Collect the last $K$ hidden states
$\mathbf{z}\leftarrow\mathcal{P}(\mathbf{G})$
// Pool generated representations into a single embedding
return $\mathbf{z}$
Algorithm 1 GIRCSE: Autoregressive Generative Embedding
Appendix E Scalability Analysis
[…]
Table 6: Computational efficiency comparison across sequence lengths (512, 1024, 2048) and generation budgets $k$. Lower is better ($\downarrow$) for both FLOPs and memory. GIRCSE without caching (red) incurs significant computational overhead due to auto-regressive processing, whereas KV caching (blue) dramatically mitigates this cost. Multipliers in parentheses show overhead relative to Causal-EOS method.
[…]
Method $\mathbf{k}$ 512 1K 2K 512 1K 2K
Causal-EOS 7.33 14.65 29.30 13.89 14.34 15.11
Bidirectional-Avg 7.33 14.65 29.30 14.03 14.45 15.24
GIRCSE (w/o cache) 1 14.67 (2.00×) 29.32 (2.00×) 58.62 (2.00×) 13.72 13.90 14.28
3 29.39 (4.01×) 58.70 (4.01×) 117.31 (4.00×) 13.74 13.92 14.29
5 44.17 (6.02×) 88.13 (6.02×) 176.04 (6.01×) 13.75 13.94 14.30
GIRCSE (w/ cache) 1 7.34 (1.00×) 14.67 (1.00×) 29.32 (1.00×) 13.73 13.91 14.32
3 7.37 (1.01×) 14.70 (1.00×) 29.35 (1.00×) 13.75 13.93 14.34
5 7.40 (1.01×) 14.73 (1.01×) 29.38 (1.00×) 13.77 13.96 14.35
Appendix F Theoretical Computational and Memory Cost Analysis
To better analyze the additional training and inference cost introduced by GIRCSE, we compare the computational and memory complexity of the proposed generative embedding framework against the conventional discriminative embedding paradigm. In the baseline discriminative case, the encoder processes an input of length $N$, leading to a per-layer cost dominated by self-attention of order $O(N^{2}d)$ and memory footprint $O(LN^{2})$, where $d$ is the embedding dimension and $L$ is the number of layers.
In the generative framework, $K$ auxiliary soft tokens are generated autoregressively. Each generation step requires an encoder forward pass over $N+j$ tokens ($j=0,\ldots,K-1$) followed by a vocabulary softmax of cost $O(d|\mathcal{V}|)$. After generation, a final encoder pass is performed over the extended sequence of length $N+K$. The total attention-dominated computation ratio with respect to the baseline is:
$R_{\text{computation}}=\frac{C_{\text{gen}}}{C_{\text{base}}}=\frac{(K+1)N^{2}+NK(K+1)+\frac{K(K+1)(2K+1)}{6}}{N^{2}},$ (9)
which simplifies to $R_{\text{computation}}\approx K+1$ when $K\ll N$. The additional softmax operations contribute $K\,O(d|\mathcal{V}|)$, which is typically small compared to the quadratic encoder cost unless $N$ is short or $|\mathcal{V}|$ is very large.
In terms of memory, peak training-time activation usage is dominated by the final encoder pass over $N+K$ tokens. Thus, the relative peak memory ratio is:
$R_{\text{memory}}=\frac{M_{\text{gen}}}{M_{\text{base}}}\approx\frac{(N+K)^{2}}{N^{2}}.$ (10)
This indicates that while the generative embedding framework incurs roughly $K$ additional encoder passes in computation, the increase in peak memory is modest, scaling quadratically with the extended sequence length $N+K$.
Appendix G Detail Implementation of Two-stage Generation Embedding
The two-stage generation embedding approach enhances representation quality by introducing an intermediate expansion step before re-encoding.
In the first stage, an auxiliary large language model (LLM) is prompted to generate a short augmentation of the input, detailed expansion prompt can be found in LABEL:tab:two-stage-prompt. The prompt instructs the model to output only the augmentation, within a fixed token budget, without explanations or additional formatting. This augmentation is designed to highlight or enrich semantic information that may be useful for downstream tasks.
In the second stage, the original instruction and text is concatenated with the generated augmentation, and the combined sequence is re-encoded into an embedding. This two-step process allows the encoder to capture a more informative and contextually aligned representation than directly embedding the raw text alone.
Table 7: LLM expansion prompt used for two-stage generation methods.
Input:
Given the INSTRUCTION and the TEXT, produce a helpful augmentation that, when concatenated to the original TEXT and embedded, is likely to improve embedding quality for the instruction’s task.
Do not explain yourself or output anything other than the augmentation.
Your answer must be written within {256} tokens.
INSTRUCTION: {instruction}
TEXT: {text}
Appendix H More Evaluation on NanoBEIR and TREC benchmarks
In this section, we provide additional evaluation results on the NanoBEIR and TREC benchmarks. These benchmarks complement the main paper by covering a wider range of retrieval tasks and domains. The results, presented in Table 8 and Table 9, highlight consistent trends with our main findings, further demonstrating the robustness and generalization ability of the proposed model compared to strong baselines.
Table 8: Performance comparison on NanoBEIR benchmark.
Dataset E5-Mistral SFR-Embedding-2 gte-Qwen2 LLM2Vec GritLM NV-Embed-v1 GIRCSE GIRCSE
Size 7B 7B 7B 7B 7B 7B 7B 7B
Vol. 1.8M 1.7M 800M 1.5M 2M 1.1M 0.2M 0.2M
Backbone Mistral Mistral QWEN2 Mistral Mistral Mistral Mistral QWEN2
ArguAna 60.13 64.33 72.37 54.41 67.72 68.00 70.18 70.88
ClimateFever 41.75 44.12 33.83 38.09 37.45 42.93 36.83 31.66
DBPedia 71.59 73.81 72.14 70.78 68.15 71.44 71.01 69.81
FEVER 95.30 95.18 80.71 96.94 94.38 96.15 94.23 95.41
FiQA 60.50 65.08 68.23 61.46 65.98 68.07 60.79 55.90
HotpotQA 86.16 91.64 82.06 88.14 91.09 91.05 88.10 88.31
MSMARCO 66.58 66.51 70.67 64.30 62.82 67.80 64.03 61.85
NFCorpus 32.36 38.99 37.67 39.19 40.39 38.54 40.75 36.57
NQ 76.64 82.20 80.99 83.81 82.91 87.11 78.57 70.01
Quora 96.28 95.48 97.44 95.64 95.76 91.59 92.73 88.76
SCIDOCS 36.00 48.18 50.76 43.72 46.32 38.11 44.27 43.97
SciFact 78.30 89.67 67.58 79.41 81.18 78.95 80.90 81.70
Touche2020 51.72 56.66 39.44 49.46 51.89 58.51 46.09 39.03
Avg. 65.64 70.14 65.68 66.57 68.16 69.10 66.81 64.14
Table 9: Performance comparison on TREC datasets. Results obtained from Weller et al. (2025) Best results per task (i.e., column) are in bold, with second best results are in underline.
Robust04 News21 Core17 Avg.
Model MAP nDCG MAP
No-Instruction IR
E5-base-v2 13.4 20.9 14.0 16.1
Contriever 19.7 22.9 15.3 19.3
MonoBERT 21.0 25.1 18.4 21.5
BM25 12.1 19.3 8.1 13.2
MonoT5-base 15.7 11.0 12.2 13.0
E5-large-v2 17.4 24.3 17.0 19.6
MonoT5-3B 27.3 16.5 18.2 20.7
Instruction-IR
TART-Contriever 14.3 21.8 13.3 16.5
INSTRUCTOR-base 17.2 22.1 15.5 18.3
E5-mistral 23.1 27.8 18.3 23.1
BGE-base 16.8 20.0 14.6 17.1
INSTRUCTOR-xl 19.7 26.1 16.8 20.9
BGE-large 17.5 22.3 15.0 18.3
GritLM-7B 28.6 24.4 20.8 24.6
TART-FLAN-T5-xl 24.6 12.8 17.0 18.1
APIs
OpenAI v3 Large 27.2 27.2 21.6 25.3
Cohere v3 English 22.3 28.3 20.6 23.7
Google Gecko 23.3 29.5 23.2 25.3
Instruct LMs
FLAN-T5-base 6.4 6.1 6.5 6.3
Llama-2-7B-chat 6.3 1.7 5.4 4.5
FLAN-T5-large 14.7 8.0 11.4 11.4
GritLM-Reranker 9.7 10.2 9.8 9.9
Mistral-7B-instruct 23.2 27.2 19.7 23.4
FollowIR-7B 24.8 29.6 20.0 24.8
End2End Generative Embedding
GIRCSE-Mistral-7B 27.9 26.8 23.0 25.9
GIRCSE-Qwen2-7B 23.9 24.1 21.0 23.0
Appendix I Full MTEB Performance
In the main paper, we reported averaged performance across task categories for clarity. Here, we provide the full per-dataset results on the MTEB benchmark, covering all tasks included in our evaluation. The detailed scores in Table 10 allow a more granular comparison across individual datasets and complement the averaged results presented in the main section.
Table 10: Full evaluation results across MTEB tasks for GIRCSE, Causal-EOS, and Bidirectional-Avg with Mistral and Qwen backbones. Best results per task (i.e., row) are in bold.
Task GIRCSE Causal-EOS Bidirectional-Avg
Mistral Qwen Mistral Qwen Mistral Qwen
MindSmallReranking 0.294 0.321 0.307 0.308 0.302 0.311
AskUbuntuDupQuestions 0.684 0.664 0.677 0.642 0.673 0.631
TwitterSemEval2015 0.769 0.747 0.760 0.732 0.786 0.707
StackExchangeClusteringP2P.v2 0.499 0.508 0.457 0.463 0.439 0.479
BiorxivClusteringP2P.v2 0.487 0.497 0.439 0.431 0.459 0.472
SICK-R 0.727 0.742 0.731 0.741 0.753 0.746
ToxicConversationsClassification 0.841 0.870 0.781 0.735 0.811 0.760
TweetSentimentExtractionClassification 0.762 0.775 0.730 0.703 0.744 0.723
TwentyNewsgroupsClustering.v2 0.611 0.627 0.598 0.582 0.624 0.569
STS15 0.845 0.833 0.835 0.817 0.839 0.815
MTOPDomainClassification 0.957 0.967 0.958 0.960 0.963 0.944
STSBenchmark 0.832 0.820 0.807 0.814 0.823 0.773
STS17 0.801 0.795 0.774 0.821 0.805 0.806
ClimateFEVERHardNegatives 0.268 0.261 0.220 0.222 0.227 0.269
HotpotQAHardNegatives 0.729 0.737 0.738 0.708 0.735 0.686
FiQA2018 0.552 0.489 0.486 0.480 0.521 0.440
CQADupstackGamingRetrieval 0.641 0.656 0.587 0.576 0.615 0.574
SprintDuplicateQuestions 0.943 0.949 0.946 0.940 0.939 0.942
ArguAna 0.699 0.692 0.694 0.574 0.703 0.662
MassiveIntentClassification 0.772 0.781 0.775 0.757 0.772 0.756
SCIDOCS 0.237 0.250 0.234 0.231 0.242 0.231
STS22.v2 0.690 0.665 0.607 0.447 0.645 0.679
STS12 0.730 0.740 0.724 0.731 0.738 0.708
STS13 0.717 0.725 0.704 0.762 0.714 0.581
MedrxivClusteringP2P.v2 0.429 0.433 0.386 0.338 0.387 0.418
MassiveScenarioClassification 0.793 0.802 0.807 0.778 0.798 0.790
STS14 0.746 0.744 0.710 0.739 0.741 0.665
ArXivHierarchicalClusteringP2P 0.645 0.630 0.639 0.646 0.634 0.639
ImdbClassification 0.960 0.962 0.955 0.906 0.959 0.949
Banking77Classification 0.855 0.861 0.849 0.838 0.853 0.823
Touche2020Retrieval.v3 0.478 0.391 0.490 0.459 0.452 0.470
SummEvalSummarization.v2 0.336 0.354 0.363 0.332 0.361 0.353
TwitterURLCorpus 0.873 0.859 0.864 0.862 0.865 0.850
AmazonCounterfactualClassification 0.887 0.918 0.893 0.731 0.896 0.867
MedrxivClusteringS2S.v2 0.423 0.428 0.421 0.409 0.423 0.403
StackExchangeClustering.v2 0.759 0.772 0.758 0.747 0.755 0.742
FEVERHardNegatives 0.853 0.857 0.786 0.702 0.809 0.835
CQADupstackUnixRetrieval 0.522 0.545 0.493 0.455 0.504 0.450
TRECCOVID 0.731 0.640 0.797 0.702 0.733 0.682
BIOSSES 0.786 0.804 0.765 0.851 0.771 0.781
ArXivHierarchicalClusteringS2S 0.648 0.634 0.644 0.645 0.645 0.657
Appendix J Full Dataset Instructions
This section provides the complete set of natural language instructions used with the datasets in our experiments. We include both the instructions for the MTEB benchmark datasets and those for the Instruction Following tasks. These instructions define the intended task for each dataset and serve as the input prompts during embedding evaluation. The full text of the instructions is listed in LABEL:tab:mteb_instruct and LABEL:tab:instruct_follow_instruct, respectively.
Table 11: Instructions for the corresponding datasets in the MTEB benchmark. We mainly follow the instructions from the GritLM paper. Note that for retrieval and reranking datasets, queries (Q) and corpus (C) documents may require different instructions, denoted as {dataset}-Q and {dataset}-C, respectively. For datasets with query instructions only (i.e., {dataset}-Q), no instructions are applied to the corpus.
Dataset Instruction
SummEvalSummarization Given a news summary, retrieve other semantically similar summaries.
ArXivHierarchicalClusteringP2P Identify the main and secondary category of Arxiv papers based on the titles and abstracts.
ArXivHierarchicalClusteringS2S Identify the main and secondary category of Arxiv papers based on the titles.
Touche2020Retrieval.v3-Q Given a question, retrieve passages that answer the question.
ClimateFEVERHardNegatives-Q Given a claim about climate change, retrieve documents that support or refute the claim.
FEVERHardNegatives-Q Given a claim, retrieve documents that support or refute the claim.
HotpotQAHardNegatives-Q Given a multi-hop question, retrieve documents that can help answer the question.
AmazonCounterfactualClassification Classify a given Amazon customer review text as either counterfactual or not-counterfactual.
AmazonPolarityClassification Classify Amazon reviews into positive or negative sentiment.
AmazonReviewsClassification Classify the given Amazon review into its appropriate rating category.
Banking77Classification Given a online banking query, find the corresponding intents.
EmotionClassification Classify the emotion expressed in the given Twitter message into one of the six emotions: anger, fear, joy, love, sadness, and surprise.
ImdbClassification Classify the sentiment expressed in the given movie review text from the IMDB dataset.
MassiveIntentClassification Given a user utterance as query, find the user intents.
MassiveScenarioClassification Given a user utterance as query, find the user scenarios.
MTOPDomainClassification Classify the intent domain of the given utterance in task-oriented conversation.
MTOPIntentClassification Classify the intent of the given utterance in task-oriented conversation.
ToxicConversationsClassification Classify the given comments as either toxic or not toxic.
TweetSentimentExtractionClassification Classify the sentiment of a given tweet as either positive, negative, or neutral.
ArxivClusteringP2P Identify the main and secondary category of Arxiv papers based on the titles and abstracts.
ArxivClusteringS2S Identify the main and secondary category of Arxiv papers based on the titles.
BiorxivClusteringP2P Identify the main category of Biorxiv papers based on the titles and abstracts.
BiorxivClusteringS2S Identify the main category of Biorxiv papers based on the titles.
MedrxivClusteringP2P Identify the main category of Medrxiv papers based on the titles and abstracts.
MedrxivClusteringS2S Identify the main category of Medrxiv papers based on the titles.
RedditClustering Identify the topic or theme of Reddit posts based on the titles.
RedditClusteringP2P Identify the topic or theme of Reddit posts based on the titles and posts.
StackExchangeClustering Identify the topic or theme of StackExchange posts based on the titles.
StackExchangeClusteringP2P Identify the topic or theme of StackExchange posts based on the given paragraphs.
TwentyNewsgroupsClustering Identify the topic or theme of the given news articles.
SprintDuplicateQuestions Retrieve duplicate questions from Sprint forum.
TwitterSemEval2015 Retrieve tweets that are semantically similar to the given tweet.
TwitterURLCorpus Retrieve tweets that are semantically similar to the given tweet.
AskUbuntuDupQuestions-Q Retrieve duplicate questions from AskUbuntu forum.
AskUbuntuDupQuestions-C Retrieve duplicate questions from AskUbuntu forum.
MindSmallReranking-Q Retrieve relevant news articles based on user browsing history.
MindSmallReranking-C Retrieve relevant news articles based on user browsing history.
SciDocsRR-Q Given a title of a scientific paper, retrieve the titles of other relevant papers.
SciDocsRR-C Given a title of a scientific paper, retrieve the titles of other relevant papers.
StackOverflowDupQuestions-Q Retrieve duplicate questions from StackOverflow forum.
StackOverflowDupQuestions-C Retrieve duplicate questions from StackOverflow forum.
ArguAna-Q Given a claim, find documents that refute the claim.
ClimateFEVER-Q Given a claim about climate change, retrieve documents that support or refute the claim.
CQADupstackRetrieval-Q Given a question, retrieve detailed question descriptions from Stackexchange that are duplicates to the given question.
DBPedia-Q Given a query, retrieve relevant entity descriptions from DBPedia.
FEVER-Q Given a claim, retrieve documents that support or refute the claim.
FiQA2018-Q Given a financial question, retrieve user replies that best answer the question.
HotpotQA-Q Given a multi-hop question, retrieve documents that can help answer the question.
MSMARCO-Q Given a web search query, retrieve relevant passages that answer the query.
NFCorpus-Q Given a question, retrieve relevant documents that best answer the question.
NQ-Q Given a question, retrieve Wikipedia passages that answer the question.
QuoraRetrieval-Q Given a question, retrieve questions that are semantically equivalent to the given question.
SCIDOCS-Q Given a scientific paper title, retrieve paper abstracts that are cited by the given paper.
SciFact-Q Given a scientific claim, retrieve documents that support or refute the claim.
Touche2020-Q Given a question, retrieve detailed and persuasive arguments that answer the question.
TRECCOVID-Q Given a query on COVID-19, retrieve documents that answer the query.
STS12 Retrieve semantically similar text.
STS13 Retrieve semantically similar text.
STS14 Retrieve semantically similar text.
STS15 Retrieve semantically similar text.
STS16 Retrieve semantically similar text.
STS17 Retrieve semantically similar text.
STS22 Retrieve semantically similar text.
BIOSSES Retrieve semantically similar text.
SICK-R Retrieve semantically similar text.
STSBenchmark Retrieve semantically similar text.
SummEval Given a news summary, retrieve other semantically similar summaries.
CQADupstackTexRetrieval-Q Represent the title of a user question to find a duplicate user question title with body from the Tex StackExchange forum.
CQADupstackTexRetrieval-C Represent the question title with body posted by a user to find a duplicate user question title from the Tex StackExchange forum.
CQADupstackWebmastersRetrieval-Q Represent the title of a user question to find a duplicate user question title with body from the Webmasters StackExchange forum.
CQADupstackWebmastersRetrieval-C Represent the question title with body posted by a user to find a duplicate user question title from the Webmasters StackExchange forum.
CQADupstackEnglishRetrieval-Q Represent the title of a user question to find a duplicate user question title with body from the English StackExchange forum.
CQADupstackEnglishRetrieval-C Represent the question title with body posted by a user to find a duplicate user question title from the English StackExchange forum.
CQADupstackGamingRetrieval-Q Represent the title of a user question to find a duplicate user question title with body from the Gaming StackExchange forum.
CQADupstackGamingRetrieval-C Represent the question title with body posted by a user to find a duplicate user question title from the Gaming StackExchange forum.
CQADupstackGisRetrieval-Q Represent the title of a user question to find a duplicate user question title with body from the Gis StackExchange forum.
CQADupstackGisRetrieval-C Represent the question title with body posted by a user to find a duplicate user question title from the Gis StackExchange forum.
CQADupstackUnixRetrieval-Q Represent the title of a user question to find a duplicate user question title with body from the Unix StackExchange forum.
CQADupstackUnixRetrieval-C Represent the question title with body posted by a user to find a duplicate user question title from the Unix StackExchange forum.
CQADupstackMathematicaRetrieval-Q Represent the title of a user question to find a duplicate user question title with body from the Mathematica StackExchange forum.
CQADupstackMathematicaRetrieval-C Represent the question title with body posted by a user to find a duplicate user question title from the Mathematica StackExchange forum.
CQADupstackStatsRetrieval-Q Represent the title of a user question to find a duplicate user question title with body from the Stats StackExchange forum.
CQADupstackStatsRetrieval-C Represent the question title with body posted by a user to find a duplicate user question title from the Stats StackExchange forum.
CQADupstackPhysicsRetrieval-Q Represent the title of a user question to find a duplicate user question title with body from the Physics StackExchange forum.
CQADupstackPhysicsRetrieval-C Represent the question title with body posted by a user to find a duplicate user question title from the Physics StackExchange forum.
CQADupstackProgrammersRetrieval-Q Represent the title of a user question to find a duplicate user question title with body from the Programmers StackExchange forum.
CQADupstackProgrammersRetrieval-C Represent the question title with body posted by a user to find a duplicate user question title from the Programmers StackExchange forum.
CQADupstackAndroidRetrieval-Q Represent the title of a user question to find a duplicate user question title with body from the Android StackExchange forum.
CQADupstackAndroidRetrieval-C Represent the question title with body posted by a user to find a duplicate user question title from the Android StackExchange forum.
CQADupstackWordpressRetrieval-Q Represent the title of a user question to find a duplicate user question title with body from the Wordpress StackExchange forum.
CQADupstackWordpressRetrieval-C Represent the question title with body posted by a user to find a duplicate user question title from the Wordpress StackExchange forum.
Table 12: Instructions for two Instruction Following datasets used in Inbedder paper.
Dataset Instruction
IntentEmotion (Intent) Represent the intent of this text.
IntentEmotion (Emotion) Represent the emotion of this text.
NYTClustering (Location) Represent the text based on where the news happen.
NYTClustering (Topic) Represent the text based on the main news category.

### v2

@@
 Our results establish generative iterative refinement as a new paradigm for representation learning.
+Our code and pre-trained models are available at https://github.com/Roytsai27/GIRCSE.
 1 Introduction
@@
 (2024).
-However, current LLM-based embeddings typically adopt an encoder-style design: embeddings are extracted in a single forward pass with contrastive learning objectives, without leveraging the generative capacity of LLMs.
+However, current LLM-based embeddings typically operate as single-pass feature extractors: embeddings are extracted in a single forward pass with contrastive learning objectives, without leveraging the generative capacity of LLMs.
 This overlooks a key strength of pretrained LLMs—their ability to reason and iteratively refine through autoregressive generation Wei et al.
@@
 (2014) captured basic semantic relationships but lacked contextual understanding.
-The introduction of transformer-based models marked a significant advancement, with BERT-based approaches like Sentence-BERT Reimers & Gurevych (2019) and SimCSE Gao et al.
+The introduction of transformer-based models marked a significant advancement, with BERT-based approaches like Sentence-BERT Reimers and Gurevych (2019) and SimCSE Gao et al.
 (2021) establishing contrastive learning as the dominant paradigm for learning sentence representations.
@@
 As summarized in Table 1, most existing approaches differ only in pooling strategies or auxiliary training techniques, while generative embeddings remain largely underexplored.
-Table 1: Comparison of LLM-based embedding models. ”Bidir.” indicates bidirectional attention, ”GP” means generated tokens pooling, ”TTS” refers to test-time scaling capability.
+Table 1: Comparison of LLM-based embedding models. “Bidir.” indicates bidirectional attention, “GP” means generated tokens pooling, “TTS” refers to test-time scaling capability.
 Method Attention Pooling LoRA Generation TTS Gen. Token Training Obj.
@@
 where $p_{\psi}$ represents the generative distribution and $\mathbf{S}_{<k}=(\mathbf{s}_{1},\ldots,\mathbf{s}_{k-1})$ are the previously generated soft tokens.
-The soft tokens are then mapped into embedding space [^1] We defer the detailed soft token generation mechanism to Section 3.2, while here we focus on the overall framework, producing $\mathbf{D}=(\mathbf{d}_{1},\ldots,\mathbf{d}_{k})\in\mathbb{R}^{K\times d}$, and is concatenated with input embeddings $\mathbf{X}$ to further feed into the Transformer encoder $f_{\theta}$:
+The soft tokens are then mapped into embedding space [^1] We defer the detailed soft token generation mechanism to Section 3.2, while here we focus on the overall framework, producing $\mathbf{D}=(\mathbf{d}_{1},\ldots,\mathbf{d}_{k})\in\mathbb{R}^{K\times d}$, and is concatenated with input embeddings $\mathbf{X}$ to further feed into the Transformer decoder $f_{\theta}$:
 $\mathbf{H}=f_{\theta}([\mathbf{X};\mathbf{D}])=(\mathbf{h}_{1}^{(L)},\mathbf{h}_{2}^{(L)},\ldots,\mathbf{h}_{N+K}^{(L)})\in\mathbb{R}^{(N+K)\times d},$ (3)
@@
 ICR combines (1) Stepwise Contrastive Loss, which supervises each generation step with contrastive loss, and (2) Iterative Refinement Regularization, which enforces progressive embedding quality improvement for each step.
+Input: Input tokens $(t_{1},\ldots,t_{N})$, generation steps $K$, embedding matrix $\mathbf{E}$, LLM decoder $f_{\theta}$, LM head parameters $W_{\phi}$ and $b_{\phi}$, pooling function $\mathcal{P}$
+Output: Final embedding $\mathbf{z}\in\mathbb{R}^{d}$
+1ex
+$\mathbf{X}\leftarrow(\mathbf{E}[t_{1}],\ldots,\mathbf{E}[t_{N}])$
+// Embed the input tokens
+$\mathbf{D}\leftarrow[\,]$
+// Initialize the list of generated embeddings
+for $k=1$ to $K$ do
+ $\mathbf{H}\leftarrow f_{\theta}([\mathbf{X};\mathbf{D}])$
+ // Forward with input and generated tokens
+ $\mathbf{h}^{\prime}_{k-1}\leftarrow\mathbf{H}[N+k-1]$
+ // Last hidden state for next-token prediction
+ $\mathbf{s}_{k}\leftarrow\mathrm{softmax}(\mathbf{W}_{\phi}\mathbf{h}^{\prime}_{k-1}+\mathbf{b}_{\phi})$
+ // Compute soft token distribution
+ $\mathbf{d}_{k}\leftarrow\sum_{i=1}^{|\mathcal{V}|}s_{k,i}\mathbf{e}_{i}$
+ // Compute soft embedding
+ $\mathbf{D}\leftarrow\mathbf{D}\|\mathbf{d}_{k}$
+ // Append soft embedding for next step
+$\mathbf{H}\leftarrow f_{\theta}([\mathbf{X};\mathbf{D}])$
+$\mathbf{G}\leftarrow(\mathbf{H}[N+1],\ldots,\mathbf{H}[N+K])$
+// Collect the last $K$ hidden states
+$\mathbf{z}\leftarrow\mathcal{P}(\mathbf{G})$
+// Pool generated representations into a single embedding
+return $\mathbf{z}$
+Algorithm 1 GIRCSE: Autoregressive Generative Embedding
 Stepwise Contrastive Loss.
@@
 Finally, the overall fine-tuning objective for generative embeddings combines the two terms: $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{contrast}}+\lambda\mathcal{L}_{\text{reg}}$, where $\lambda$ is a hyperparameter that balances contrastive alignment and refinement regularization.
+We summarize the overall training procedure in Algorithm 2.
+During each training iteration, we sample a mini-batch of queries and documents.
+For every input sequence, the model performs autoregressive soft token generation for $K$ steps to obtain the sequence of embeddings.
+Finally, the model parameters $\psi$ are updated end-to-end via gradient descent to minimize the combined objective $\mathcal{L}_{\text{total}}$.
 4 Experiment
@@
 (2023b), BGE-Large Li et al.
-(2023a), and UAE-Large Li & Li (2024).
+(2023a), and UAE-Large Li and Li (2024).
 (2) LLM-based methods are instruction-tuned LLM embeddings, including LLM2Vec BehnamGhader et al.
@@
 (2023b).
-(3) Generative embeddings cover (i) two-stage approaches that expand text with an auxiliary LLM before re-encoding (see Appendix G for detail) and (ii) the end-to-end generative model Inbedder Peng et al.
+(3) Generative embeddings cover (i) two-stage approaches that expand text with an auxiliary LLM before re-encoding (see Appendix F for detail) and (ii) the end-to-end generative model Inbedder Peng et al.
 (2024).
@@
 For detailed performance of each MTEB dataset, please refer to Appendix I.
-MTEB (English, v2) Instruct Following
-Overall
-Rank
+MTEB (English, v2) Instruct Following Overall Rank
 Task Size Vol. Backbone Retr. Rerank. Clust. PairClass. Class. STS Summ. Avg. (Rank) IntEmo NYT Avg. (Rank)
@@
 These results highlight autoregressive generation as a powerful mechanism for embedding optimization and open new directions for scalable, semantically rich representations.
+Acknowledgments
+This material is based upon work supported by National Science and Technology Council, ROC under grant number 114-2221-E-002-134-MY3 and by Taiwan Centers of Excellence (TCE)
 Appendix A Use-of-LLMs
@@
 During inference, we increase $K$ to 20 to enable longer generations and improved embedding refinement, while mitigating the computational overhead using KV-cache techniques.
-Appendix D Pseudo Code for GIRCSE
-Algorithm 1 provides a pseudo-code summary of GIRCSE.
-We include this block to clarify the core computation steps and facilitate reproducibility.
-The algorithm specifies how input tokens are embedded, how soft embeddings are autoregressively generated across multiple steps, and how the final representation is obtained through pooling.
-While implementation details may vary, this summary highlights the essential components needed to reproduce our method.
 Input: Input tokens $(t_{1},\ldots,t_{N})$, generation steps $K$, embedding matrix $\mathbf{E}$, LLM decoder $f_{\theta}$, LM head parameters $W_{\phi}$ and $b_{\phi}$, pooling function $\mathcal{P}$
@@
 Algorithm 2 GIRCSE: Autoregressive Generative Embedding
-Appendix E Scalability Analysis
+Appendix D Scalability Analysis
 […]
@@
 5 7.40 (1.01×) 14.73 (1.01×) 29.38 (1.00×) 13.77 13.96 14.35
-Appendix F Theoretical Computational and Memory Cost Analysis
+Appendix E Theoretical Computational and Memory Cost Analysis
 To better analyze the additional training and inference cost introduced by GIRCSE, we compare the computational and memory complexity of the proposed generative embedding framework against the conventional discriminative embedding paradigm.
@@
 This indicates that while the generative embedding framework incurs roughly $K$ additional encoder passes in computation, the increase in peak memory is modest, scaling quadratically with the extended sequence length $N+K$.
-Appendix G Detail Implementation of Two-stage Generation Embedding
+Appendix F Detail Implementation of Two-stage Generation Embedding
 The two-stage generation embedding approach enhances representation quality by introducing an intermediate expansion step before re-encoding.
@@
 TEXT: {text}
+Appendix G Training Stability Analysis
+A potential concern when optimizing models that involve the generation of soft tokens is the risk of gradient instability.
+To empirically validate the stability of our proposed model, GIRCSE, we monitor its training dynamics.
+We present the training loss curve in Figure 4 and the L2 norm of the gradients in Figure 5.
+As shown in the figures, the training loss (Figure 4) demonstrates a smooth and consistent decrease, indicating stable convergence.
+Furthermore, the gradient norm (Figure 5) converges normally throughout the training process.
+Figure 4: Training loss curve of GIRCSE.
+Figure 5: Training gradient norm (L2) of GIRCSE, plotted with the top 2% of outliers removed for clarity.
 Appendix H More Evaluation on NanoBEIR and TREC benchmarks